OpenStudy (anonymous):

List the integers 1 to 10 inclusive that are even and a multiple of 3.

OpenStudy (austinl):

3,6?

OpenStudy (anonymous):

6

OpenStudy (austinl):

3,6,9? 3*1=3 3*2=6 3*3=9

OpenStudy (anonymous):

6 is the only even integer multiple of 3 from 1to 10

OpenStudy (austinl):

Oopsies, missed the "even" part. Good catch. Medal for you sir.

OpenStudy (anonymous):

@austinL Thank you